A health specialist wants to determine the average number of hours a person exercise

in a day during the quarantine period. She found out that the mean number of hours a

person exercise in a day during the quarantine period is 80 minutes. A random sample of

29 persons were surveyed and found that their mean is 65 minutes and a standard

deviation of 10 minutes. Test the hypothesis at 2% level of significance and assume that

the population is normally distributed.


1
Expert's answer
2022-05-17T11:57:53-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu=80"

"H_a:\mu\not=80"

This corresponds to a two-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha = 0.02," "df=n-1=28" degrees of freedom, and the critical value for a two-tailed test is "t_c = 2.46714."The rejection region for this two-tailed test is "R = \{t: |t| > 2.46714\}."

The t-statistic is computed as follows:


"t=\dfrac{65-80}{10/\sqrt{29}}=-8.0777"

Since it is observed that "|t| = 8.0777> 2.467114=t_c," it is then concluded that the null hypothesis is rejected.

Using the P-value approach: The p-value for two-tailed "df=28" degrees of freedom, "t=-8.0777" is "p\approx 0," and since "p=0<0.02=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu" is different than 80, at the "\alpha = 0.02" significance level.